# Can Large Multimodal Models Understand Agricultural Scenes? Benchmarking with AgroMind

### v2

@@ Abstract
Large Multimodal Models (LMMs) has demonstrated capabilities across various domains, but comprehensive benchmarks for agricultural remote sensing (RS) remain scarce. Existing benchmarks designed for agricultural RS scenarios exhibit notable limitations, primarily in terms of insufficient scene diversity in the dataset and oversimplified task design. To bridge this gap, we introduce AgroMind, a comprehensive agricultural remote sensing benchmark covering four task dimensions: spatial perception, object understanding, scene understanding, and scene reasoning, with a total of 13 task types, ranging from crop identification and health monitoring to environmental analysis. We curate a high-quality evaluation set by integrating eight public datasets and one private farmland plot dataset, containing 27,247 QA pairs and 19,615 images. The pipeline begins with multi-source data pre-processing, including collection, format standardization, and annotation refinement. We then generate a diverse set of agriculturally relevant questions through the systematic definition of tasks. Finally, we employ LMMs for inference, generating responses, and performing detailed examinations. We evaluated 20 open-source LMMs and 4 closed-source models on AgroMind. Experiments reveal significant performance gaps, particularly in spatial reasoning and fine-grained recognition, it is notable that human performance lags behind several leading LMMs. By establishing a standardized evaluation framework for agricultural RS, AgroMind reveals the limitations of LMMs in domain knowledge and highlights critical challenges for future work. Data and code can be accessed at https://rssysu.github.io/AgroMind/.
@@ 1 Introduction
AgroMind (ours) 13 types 27K
To address these gaps, we introduce AgroMind, a multi-scenario, multi-task benchmark designed to comprehensively evaluate LMMs in the agricultural RS environment. We integrate 8 public datasets and 1 private farmland plot dataset to create a multi-scene dataset supporting various critical agricultural applications. AgroMind covers four core task dimensions: spatial perception, object understanding, scene understanding, and scene reasoning, with a total of 13 task types, ranging from crop identification and health monitoring to environmental analysis. AgroMind includes both basic tasks from previous benchmarks as well as reasoning tasks in agricultural real-world scenarios.
@@
$\bullet$ We construct a large-scale multi-modal and multi-scene agricultural remote sensing dataset, which includes 8 public datasets and 1 private farmland plot dataset, containing 19,615 images and 27,247 QA pairs. The dataset spans critical applications such as pest and disease monitoring, plot analysis, multiple crop recognition, as well as anomaly detection in agricultural scenarios.
@@ 3.1 Dataset statistics
This section presents fundamental statistical characteristics of the AgroMind dataset. The dataset is constructed by integrating eight publicly available RS datasets and one proprietary farmland plot dataset, specifically including OAM-TCD veitchmichaelis2024oamtcdgloballydiversedataset , OilPalmUAV zheng2021growing , AVC Chiu_2020_CVPR , PhenoBench weyler2024pami , Maize Tassel Identification Dataset (Iflytek AI Competition 2024), IP102 Wu2019Insect , ACFR bargoti2016deep , 2018 AI Challenger Dataset, and 7-Provinces Ground Dataset. Through a selection process, we retained 19,615 images and generated 27,247 question-answer (QA) pairs from these visual data, ensuring comprehensive diversity across agricultural scenarios. The proportional contributions from different source datasets are quantitatively illustrated in Fig. 3 (a). More details can be found in Appendix B.1.
Statistics for different sensors. The AgroMind dataset was compiled from three sensor types: unmanned aerial vehicles (UAVs), satellite platforms, and cameras. Fig. 3 (b) presents the exact numerical distribution of images and corresponding QA pairs for each sensor type, revealing a QA pair ratio of approximately 9:11:7 for ground-based cameras, satellite platforms, and UAVs, respectively. Notably, while satellite and ground imagery constitute the majority of the dataset, UAV-derived annotations still encompass 7,000 QA pairs, ensuring sufficient coverage to validate model robustness in multi-perspective agricultural tasks.
Statistics for different scenes. Fig. 3 (c) demonstrates the dataset’s comprehensive coverage of agricultural scenarios spanning five critical categories: Anomaly Detection, Crop Monitoring, Pest Identification, Plot Delineation, and Tree Analysis. The Crop Monitoring category contains approximately 2,000 high-quality images paired with 7,000 annotated QA pairs, forming the largest data subset. The Pest Identification and Tree Analysis categories consist of 5,000 images (with 6,000 QA pairs) and 6,000 images (with 6,000 QA pairs), respectively. Notably, the inclusion of specialized scenarios such as Anomaly Detection and Plot Delineation demonstrates the dataset’s capacity to address emerging challenges in precision agriculture.
@@ 3.2 Dataset preprocessing
Customized processing protocols were applied to heterogeneous data sources: (1) TIFF-format remote sensing images were converted to PNG/JPG formats to ensure model compatibility and prevent parsing failures; (2) High-resolution geospatial images underwent randomized parcel-boundary cropping to generate multi-scale samples simulating spatial query variations; (3) A manual screening step was performed on augmented images to remove samples exhibiting artifacts such as content overlap or excessive flipping that could distort semantic information;(4) Instance-level statistical analysis and bounding box annotation were performed using original labels and masks. The resulting standardized dataset contains hierarchical annotations across three levels: pixel-level (segmentation masks), instance-level (object detection boxes), and parcel-level (agricultural boundaries). Original image resolutions were preserved (300×300 to 4,500×4,500 pixels) to maintain data integrity. Moreover, we applied a logic enhancement strategy in the QA pair design, artificially predefining the hierarchical reasoning process instead of directly using the raw annotations as the answers of QA pairs. Complete collection procedures and statistical details are documented in the Appendix B.
@@ 4 AgroMind benchmark
We introduce AgroMind, a comprehensive multimodal benchmark designed to assess the capabilities of LMMs in agricultural scenarios, containing 27,247 image-question-answer pairs. The AgroMind is characterized by the following features: (1) Broad task coverage. AgroMind covers four categories of tasks (Fig. 4(a)): spatial perception, object understanding, scene understanding and reasoning, and a total of 13 task types, covering typical visual language requirements in agricultural production. (2) versatile evaluation. It supports a wide range of question formats such as judgment, counting, image choice, text choice, and open-ended questions, allowing for comprehensive evaluation from basic perception to high-level reasoning (Fig. 4(b)). (3) Domain-specific focus. As visualized in the word clouds, the benchmark emphasizes agricultural-specific terms such as cultivated, land, pest, healthy and anomaly, demonstrating its alignment of the domain (Fig. 4(c)).
@@ 5.1 Evaluation setups
Human 23.18 19.31 46.25 58.35 47.79 49.58 35.25 30.72 37.75 15.71 25.06 32.04 18.77 34.72
Random 17.38 10.83 11.00 29.47 25.80 11.50 23.96 7.69 22.49 8.54 18.24 20.27 26.56 19.25
GPT-4o openai2024gpt4o 25.71 13.50 20.50 81.62 49.70 45.42 36.81 33.34 45.75 34.00 30.05 58.76 27.39 42.98
Gemini-1.5-Flash team2023gemini 25.00 17.50 28.75 75.42 48.05 65.17 34.64 29.07 39.11 33.06 23.92 60.82 28.06 41.21
Gemini-1.5-Pro team2023gemini 31.69 28.33 30.00 73.95 53.50 30.00 35.98 27.27 40.18 17.55 19.08 68.60 28.33 38.99
Claude-3.5-Sonnet anthropic2024claude3 41.12 17.00 29.00 50.63 45.35 29.50 29.75 22.55 34.64 26.54 14.91 40.09 28.77 33.33
DeepSeek-VL2-tiny deepseekvl2 17.86 12.00 16.25 57.88 25.75 45.92 27.88 27.58 12.55 14.02 10.75 37.00 26.98 27.51
DeepSeek-VL2-small deepseekvl2 19.88 13.33 21.25 59.91 31.15 37.92 28.36 27.67 23.50 14.35 16.35 31.96 7.20 27.84
TinyLLaVA TinyLLaVA 35.02 12.00 30.00 63.08 22.20 40.25 26.98 28.60 31.58 7.99 15.90 38.95 10.12 29.88
InternVL2-2B InternVL2 40.03 15.00 24.75 40.28 27.30 47.50 25.65 25.86 25.73 12.84 28.84 36.88 19.65 28.49
InternVL2-4B InternVL2 28.59 12.50 28.75 48.18 45.10 13.00 30.54 25.86 39.57 16.02 24.15 41.01 25.52 31.46
XComposer2-4KHD XComposer2 32.08 11.83 40.00 46.18 37.20 35.75 32.71 27.67 27.11 10.23 38.08 43.53 38.47 32.31
LLaVA-NeXT-7B-Mistral LLaVA-NeXT 41.12 12.00 49.25 43.45 42.05 61.00 20.16 28.55 39.93 7.87 11.88 30.93 15.86 31.07
LLaVA-NeXT-7B-Vicuna LLaVA-NeXT 7.14 12.00 48.75 47.68 29.75 25.83 18.71 27.76 22.15 9.91 10.22 32.42 20.65 24.61
InstructBLIP-Vicuna-7B InstrucBLIP 10.84 11.83 14.75 38.09 24.10 16.42 18.47 15.15 15.51 4.32 17.18 22.11 8.08 17.70
LLaVA-NeXT-Interleave-7B LLaVA-NeXT-Interleave 9.42 8.50 25.25 29.75 18.55 3.50 21.00 16.35 19.31 10.84 13.47 23.71 19.40 17.83
Mantis-LLaMA3-SigLIP Mantis 20.75 10.83 28.75 63.47 36.70 16.08 28.24 31.32 26.19 12.68 13.55 42.73 36.59 31.60
Mantis-Idefics2 Mantis 16.34 11.67 33.75 62.36 46.25 39.92 21.97 27.84 23.65 9.91 14.31 47.65 36.64 31.84
LLaVA-NeXT-8B LLaVA-NeXT 45.42 12.17 44.50 51.99 34.75 39.92 25.59 29.63 35.10 8.32 13.10 37.57 25.27 32.07
InternVL2-8B InternVL2 39.32 11.17 23.50 52.21 41.85 38.25 33.19 28.49 39.50 18.96 16.81 48.34 31.10 34.88
Idefics-2-8b Idefics2 32.57 13.83 33.50 60.30 36.30 42.58 23.72 27.64 23.59 9.74 14.91 44.10 23.69 30.85
LLaVA-NeXT-13B LLaVA-NeXT 38.56 12.00 47.25 55.13 26.70 34.67 20.64 29.95 38.56 5.87 11.13 32.53 31.77 31.43
InternVL2-26B InternVL2 21.46 13.50 18.25 69.50 43.90 28.33 32.47 28.37 45.13 17.57 12.72 57.62 25.77 35.84
LLaVA-NeXT-34B LLaVA-NeXT 48.15 9.33 54.50 57.74 43.00 37.08 22.51 29.10 39.81 13.62 12.72 42.15 27.60 34.93
GeoChat kuckreja2023geochat 32.79 8.33 30.00 44.26 27.00 38.33 20.73 24.05 16.56 8.57 18.32 41.86 27.50 26.15
GeoLLaVA-8K wang2025geollava 9.29 11.67 12.50 27.45 21.00 10.83 18.29 8.21 19.63 7.76 9.92 29.07 7.92 15.24
@@ 5.2 Main results
Overall challenge presented in AgroMind. As indicated in Table 2, AgroMind poses significant challenges to current SOTA LMMs. Although some models, such as GPT-4o, have surpassed human performance, 18 out of the 24 evaluated models still underperform compared to humans. Both close-sourced and open-sourced models fall short of consistently outperforming humans. The challenging nature of our benchmark indicates that current LMMs’ strong performance on the standard tasks are not generalized to the multi-modal and multi-scene agricultural scenarios. Additionally, domain-specific models like GeoChat kuckreja2023geochat and GeoLLaVA-8K wang2025geollava , while effective in general remote sensing tasks, demonstrate limitations and reduced generalization in complex agricultural scenarios.
LMMs’ performances across task dimensions. Table 2 summarizes the performance of various LMMs across 13 tasks. Most LMMs perform poorly on Spatial Perception tasks such as SR, indicating difficulties in analyzing the spatial distribution of crops or farmland. While LMMs achieve relatively high results on Object Understanding tasks like OC and PDD, their performance drops significantly on GSR, which is crucial for crop health monitoring. For Scene Understanding, most LMMs underperform compared to humans, indicating that they tend to reproduce answers seen in their training data rather than understand the agricultural scene itself. Interestingly, LMMs excel at Scene Reasoning tasks like CTR and PL by using their broad knowledge base to give expert answers. However, they perform much worse than humans on VPR, which requires visual reasoning, indicating that LMMs still need significant improvement for vision-intensive agricultural tasks.
@@
Human 78.36 46.44 15.36 34.72
Random 19.28 26.41 10.53 19.25
GPT-4o 73.66 56.56 24.56 42.98
Gemini-1.5-Flash 79.62 52.01 23.78 41.21
Gemini-1.5-Pro 66.06 53.97 23.81 38.99
Claude-3.5-Sonnet 53.40 45.73 18.07 33.33
DeepSeek-VL2-tiny 73.53 31.86 14.44 27.51
DeepSeek-VL2-small 60.98 35.24 14.55 27.84
TinyLLaVA 67.11 36.36 16.61 29.88
InternVL2-2B 62.55 31.94 18.60 28.49
InternVL2-4B 53.57 44.77 14.97 31.46
XComposer2-4KHD 59.91 41.61 18.39 32.31
LLaVA-NeXT-7B-Mistral 69.32 37.70 17.47 31.07
LLaVA-NeXT-7B-Vicuna 55.40 30.11 13.50 24.61
InstructBLIP-Vicuna-7B 34.72 23.06 9.45 17.70
LLaVA-NeXT-Interleave-7B 25.79 22.83 11.72 17.83
Mantis-LLaMA3-SigLIP 58.47 44.53 14.51 31.60
Mantis-Idefics2 65.40 43.18 14.88 31.84
LLaVA-NeXT-8B 60.38 40.64 18.66 32.07
InternVL2-8B 62.64 45.74 19.50 34.88
Idefics-2-8b 61.45 39.41 17.00 30.85
LLaVA-NeXT-13B 59.91 42.30 15.90 31.43
InternVL2-26B 64.55 50.66 16.68 35.84
LLaVA-NeXT-34B 61.28 44.69 20.84 34.93
GeoChat 56.56 26.35 20.95 26.15
GeoLLaVA-8K 19.00 24.87 7.76 15.24
@@
Is model size directly linked to performance? Fig. 7 illustrates the performance of a series of models with varying parameter scales. While larger models generally tend to achieve better performance, the overall performance is more influenced by the model’s architecture rather than its parameter size. For instance, the LLaVA-NeXT model with 34B parameters performs worse than the InternVL2 model with 26B parameters. Furthermore, when the parameter size of LLaVA-NeXT increases from 8B to 13B, its accuracy even drops by 0.64%.
Does GPT-4o exceed human performance? As shown in Table 2, GPT-4o’s overall accuracy surpasses human performance by 8.26%. We attribute this to the fact that our human participants are primarily students, who may lack the specialized knowledge included in GPT-4o’s extensive knowledge base and GPT-4o is the most SOTA model, which demonstrates a substantial lead over all other evaluated models. However, despite outperforming humans in most tasks, GPT-4o falls far behind in tasks such as BD and GSR, which exposes its limitations in spatial perception and understanding of growth states in agriculture. More results can be found in Appendix C.5.

### v3

@@ Abstract
Large Multimodal Models (LMMs) has demonstrated capabilities across various domains, but comprehensive benchmarks for agricultural remote sensing (RS) remain scarce. Existing benchmarks designed for agricultural RS scenarios exhibit notable limitations, primarily in terms of insufficient scene diversity in the dataset and oversimplified task design. To bridge this gap, we introduce AgroMind, a comprehensive agricultural remote sensing benchmark covering four task dimensions: spatial perception, object understanding, scene understanding, and scene reasoning, with a total of 13 task types, ranging from crop identification and health monitoring to environmental analysis. We curate a high-quality evaluation set by integrating nine public datasets and one private global parcel dataset, containing 28,482 QA pairs and 20,850 images. The pipeline begins with multi-source data pre-processing, including collection, format standardization, and annotation refinement. We then generate a diverse set of agriculturally relevant questions through the systematic definition of tasks. Finally, we employ LMMs for inference, generating responses, and performing detailed examinations. We evaluated 20 open-source LMMs and 4 closed-source models on AgroMind. Experiments reveal significant performance gaps, particularly in spatial reasoning and fine-grained recognition, it is notable that human performance lags behind several leading LMMs. By establishing a standardized evaluation framework for agricultural RS, AgroMind reveals the limitations of LMMs in domain knowledge and highlights critical challenges for future work. Data and code can be accessed at https://rssysu.github.io/AgroMind/.
@@ 1 Introduction
AgroMind (ours) 13 types 28K
To address these gaps, we introduce AgroMind, a multi-scenario, multi-task benchmark designed to comprehensively evaluate LMMs in the agricultural RS environment. We integrate 9 public datasets and 1 private global parcel dataset to create a multi-scene dataset supporting various critical agricultural applications. AgroMind covers four core task dimensions: spatial perception, object understanding, scene understanding, and scene reasoning, with a total of 13 task types, ranging from crop identification and health monitoring to environmental analysis. AgroMind includes both basic tasks from previous benchmarks as well as reasoning tasks in agricultural real-world scenarios.
@@
$\bullet$ We construct a large-scale multi-modal and multi-scene agricultural remote sensing dataset, which includes 9 public datasets and 1 private global parcel dataset, containing 20,850 images and 28,482 QA pairs. The dataset spans critical applications such as pest and disease monitoring, parcel analysis, multiple crop recognition, as well as anomaly detection in agricultural scenarios.
@@ 3.1 Dataset statistics
This section presents fundamental statistical characteristics of the AgroMind dataset. The dataset is constructed by integrating nine publicly available RS datasets and one proprietary global parcel dataset, specifically including OAM-TCD veitchmichaelis2024oamtcdgloballydiversedataset , OilPalmUAV zheng2021growing , AVC Chiu_2020_CVPR , PhenoBench weyler2024pami , Maize Tassel Identification Dataset (Iflytek AI Competition 2024), IP102 Wu2019Insect , ACFR bargoti2016deep , 2018 AI Challenge Dataset, CropHarvesttseng2021cropharvest , and Global Parcel Dataset. Through a selection process, we retained 20,850 images and generated 28,482 question-answer (QA) pairs from these visual data, ensuring comprehensive diversity across agricultural scenarios. The proportional of different source datasets are quantitatively illustrated in Fig. 3 (a). More details can be found in Appendix B.1.
Statistics for different sensors. The AgroMind dataset was compiled from three sensor types: unmanned aerial vehicles (UAVs), satellite platforms, and cameras. Fig. 3 (b) presents the exact numerical distribution of images and corresponding QA pairs for each sensor type, revealing a QA pair ratio of approximately 9:12:7 for ground-based cameras, satellite platforms, and UAVs, respectively. Notably, while satellite and ground imagery constitute the majority of the dataset, UAV-derived annotations still encompass 7,000 QA pairs, ensuring sufficient coverage to validate model robustness in multi-perspective agricultural tasks.
Statistics for different scenes. Fig. 3 (c) demonstrates the comprehensive coverage of agricultural scenarios spanning five critical categories: Anomaly Detection, Crop Monitoring, Pest Identification, Parcel Delineation, and Tree Analysis. The Crop Monitoring category contains approximately 3,500 high-quality images paired with 8,000 annotated QA pairs, forming the largest data subset. The Pest Identification and Tree Analysis categories consist of 5,000 images (with 6,000 QA pairs) and 6,000 images (with 6,000 QA pairs), respectively. Notably, the inclusion of specialized scenarios such as anomaly detection and parcel delineation demonstrates the capacity to address emerging challenges in precision agriculture.
@@ 3.2 Dataset preprocessing
Customized processing protocols were applied to heterogeneous data sources: (1) TIFF-format remote sensing images were processed based on content: Colorful images recognizable to the human eye were converted to PNG/JPG formats to ensure model compatibility and prevent parsing failures; while images containing primarily non-visible spectral bands had all bands converted into grayscale blocks that were concatenated, preserving the original spectral information through grayscale representation; (2) High-resolution geospatial images underwent randomized parcel-boundary cropping to generate multi-scale samples simulating spatial query variations; (3) A manual screening step was performed on augmented images to remove samples exhibiting artifacts such as content overlap or excessive flipping that could distort semantic information;(4) Instance-level statistical analysis and bounding box annotation were performed using original labels and masks. The resulting standardized dataset contains hierarchical annotations across three levels: pixel-level (segmentation masks), instance-level (object detection boxes), and parcel-level (agricultural boundaries). Original image resolutions were preserved (300×300 to 4,500×4,500 pixels) to maintain data integrity. Moreover, we applied a logic enhancement strategy in the QA pair design, artificially predefining the hierarchical reasoning process instead of directly using the raw annotations as the answers of QA pairs. Complete collection procedures and statistical details are documented in the Appendix B.
@@ 4 AgroMind benchmark
We introduce AgroMind, a comprehensive multimodal benchmark designed to assess the capabilities of LMMs in agricultural scenarios, containing 28,482 image-question-answer pairs. The AgroMind is characterized by the following features: (1) Broad task coverage. AgroMind covers four categories of tasks (Fig. 4(a)): spatial perception, object understanding, scene understanding and reasoning, and a total of 13 task types, covering typical visual language requirements in agricultural production. (2) versatile evaluation. It supports a wide range of question formats such as judgment, counting, image choice, text choice, and open-ended questions, allowing for comprehensive evaluation from basic perception to high-level reasoning (Fig. 4(b)). (3) Domain-specific focus. As visualized in the word clouds, the benchmark emphasizes agricultural-specific terms such as cultivated, land, pest, healthy and anomaly, demonstrating its alignment of the domain (Fig. 4(c)).
@@ 5.1 Evaluation setups
Human 19.15 15.00 40.83 57.60 46.17 47.22 35.49 28.88 39.46 12.33 25.62 34.07 13.50 33.15
Random 18.03 17.97 16.48 35.65 25.65 10.00 24.39 4.99 21.20 9.19 19.87 21.70 26.96 19.24
GPT-4o openai2024gpt4o 41.38 35.16 33.52 78.55 66.09 55.83 43.90 23.75 46.20 16.10 27.15 69.81 30.90 43.14
Gemini-1.5-Flash team2023gemini 29.56 39.06 24.73 77.68 62.17 55.83 33.54 23.75 43.04 22.88 29.14 68.87 27.90 40.92
Gemini-1.5-Pro team2023gemini 27.59 23.44 24.73 77.39 52.61 45.83 39.02 32.30 47.78 19.92 34.44 66.04 30.04 41.06
Claude-3.5-Sonnet anthropic2024claude3 24.14 17.19 40.11 57.97 50.00 57.50 44.51 24.47 35.44 20.34 33.11 50.94 36.05 37.11
DeepSeek-VL2-tiny deepseekvl2 16.75 24.22 19.78 59.42 25.22 55.83 29.88 21.38 14.87 17.37 15.89 44.34 21.89 27.51
DeepSeek-VL2-small deepseekvl2 16.26 20.31 24.18 61.45 28.70 65.83 31.71 19.71 20.57 19.07 19.87 38.68 8.58 28.08
TinyLLaVA TinyLLaVA 29.56 17.97 24.73 68.41 23.48 30.00 24.39 23.99 31.65 3.39 13.91 45.28 9.44 28.01
InternVL2-2B InternVL2 53.69 19.53 17.03 41.16 30.43 42.50 26.22 21.62 29.11 14.83 27.15 39.62 14.16 28.40
InternVL2-4B InternVL2 29.06 23.44 25.27 48.70 44.78 19.17 30.49 19.00 46.84 19.49 30.46 50.94 12.88 31.15
XComposer2-4KHD XComposer2 29.06 27.34 32.97 51.88 40.00 44.17 34.15 22.33 27.53 11.86 37.75 43.40 38.63 33.02
LLaVA-NeXT-7B-Mistral LLaVA-NeXT 33.50 22.66 40.11 44.64 40.00 36.67 21.95 24.70 36.71 8.05 15.23 39.62 11.59 29.17
LLaVA-NeXT-7B-Vicuna LLaVA-NeXT 4.43 22.66 40.11 52.75 30.43 22.50 20.12 20.90 21.84 14.41 11.26 40.57 24.89 25.82
InstructBLIP-Vicuna-7B InstrucBLIP 9.85 16.41 20.88 33.33 21.74 10.83 19.51 10.93 21.52 4.66 20.53 20.75 11.16 17.39
LLaVA-NeXT-Interleave-7B LLaVA-NeXT-Interleave 12.32 15.62 22.53 32.46 20.87 4.17 21.95 10.69 20.25 13.98 18.54 29.25 21.89 19.01
Mantis-LLaMA3-SigLIP Mantis 23.65 21.09 25.27 65.51 32.17 12.50 29.88 26.84 33.54 12.29 21.19 46.23 30.04 31.18
Mantis-Idefics2 Mantis 16.75 20.31 31.32 63.77 44.35 37.50 22.56 24.23 20.89 8.47 13.91 56.60 30.90 30.41
LLaVA-NeXT-8B LLaVA-NeXT 35.47 16.41 41.76 57.68 28.70 30.00 19.51 26.84 34.81 10.59 17.22 37.74 33.48 31.53
InternVL2-8B InternVL2 24.63 23.44 19.23 72.17 43.48 24.17 32.93 23.75 54.11 22.46 18.54 61.32 27.90 36.30
Idefics-2-8b Idefics2 21.67 14.84 26.37 58.84 33.91 32.50 21.95 23.04 19.94 13.98 22.52 49.06 9.44 27.09
LLaVA-NeXT-13B LLaVA-NeXT 45.32 19.53 36.81 51.88 33.04 31.67 23.17 22.09 34.81 5.51 21.85 42.45 26.18 30.69
InternVL2-26B InternVL2 46.31 24.22 19.78 50.72 42.61 55.00 39.02 24.47 44.30 33.05 24.50 56.60 31.76 37.25
LLaVA-NeXT-34B LLaVA-NeXT 49.26 16.41 39.56 59.42 40.00 30.83 25.61 27.32 43.35 17.80 19.87 45.28 28.33 35.52
GeoChat kuckreja2023geochat 14.78 19.53 23.63 48.70 34.78 30.83 16.46 19.48 20.57 8.47 29.80 51.89 24.89 25.93
GeoLLaVA-8K wang2025geollava 9.85 25.78 17.58 34.78 15.22 7.50 17.68 4.75 16.14 7.63 14.57 34.91 8.58 15.73
@@ 5.2 Main results
Overall challenge presented in AgroMind. As indicated in Table 2, AgroMind poses significant challenges to current SOTA LMMs. Although some models, such as GPT-4o, have surpassed human performance, 17 out of the 24 evaluated models still underperform compared to humans. Both close-sourced and open-sourced models fall short of consistently outperforming humans on all tasks. The challenging nature of our benchmark indicates that current LMMs’ strong performance on the standard tasks are not generalized to the multi-modal and multi-scene agricultural scenarios. Additionally, domain-specific models like GeoChat kuckreja2023geochat and GeoLLaVA-8K wang2025geollava , while effective in general remote sensing tasks, demonstrate limitations in complex agricultural scenarios.
LMMs’ performances across task dimensions. Table 2 summarizes the performance of various LMMs across 13 tasks. Most LMMs perform poorly on Spatial Perception tasks such as BD, indicating difficulties in analyzing the spatial distribution of crops or farmland. While LMMs achieve relatively high results on Object Understanding tasks like OC and PDD, their performance drops significantly on GSR, which is crucial for crop health monitoring. For Scene Understanding, most LMMs underperform compared to humans, indicating that they tend to reproduce answers seen in their training data rather than understand the agricultural scene itself. Interestingly, LMMs excel at Scene Reasoning tasks like CTR and PL by using their broad knowledge base to give expert answers. However, they perform much worse than humans on VPR, which requires visual reasoning, indicating that LMMs still need significant improvement for vision-intensive agricultural tasks.
@@
Human 78.01 45.06 13.36 33.15
Random 23.21 25.84 11.41 19.24
GPT-4o openai2024gpt4o 78.16 55.77 24.29 43.14
Gemini-1.5-Flash team2023gemini 77.47 53.68 21.62 40.92
Gemini-1.5-Pro team2023gemini 75.09 53.68 22.44 41.06
Claude-3.5-Sonnet anthropic2024claude3 66.21 51.09 18.35 37.11
DeepSeek-VL2-tiny deepseekvl2 72.35 31.02 14.64 27.51
DeepSeek-VL2-small deepseekvl2 70.31 33.03 14.49 28.08
TinyLLaVA TinyLLaVA 60.41 35.12 14.64 28.01
InternVL2-2B InternVL2 55.63 31.10 20.06 28.40
InternVL2-4B InternVL2 54.27 42.06 16.42 31.15
XComposer2-4KHD XComposer2 61.09 40.47 20.28 33.02
LLaVA-NeXT-7B-Mistral LLaVA-NeXT 57.68 37.37 15.68 29.17
LLaVA-NeXT-7B-Vicuna LLaVA-NeXT 49.83 32.78 14.41 25.82
InstructBLIP-Vicuna-7B InstrucBLIP 34.47 20.57 10.85 17.39
LLaVA-NeXT-Interleave-7B LLaVA-NeXT-Interleave 28.67 23.33 13.08 19.01
Mantis-LLaMA3-SigLIP Mantis 54.61 43.06 15.53 31.18
Mantis-Idefics2 Mantis 61.43 42.14 13.22 30.41
LLaVA-NeXT-8B LLaVA-NeXT 57.34 42.06 16.57 31.53
InternVL2-8B InternVL2 64.51 49.41 18.50 36.30
Idefics-2-8b Idefics2 57.68 33.86 14.41 27.09
LLaVA-NeXT-13B LLaVA-NeXT 56.31 39.30 17.46 30.69
InternVL2-26B InternVL2 67.58 46.15 22.73 37.25
LLaVA-NeXT-34B LLaVA-NeXT 59.04 46.40 20.73 35.52
GeoChat kuckreja2023geochat 51.54 33.28 13.82 25.93
GeoLLaVA-8K wang2025geollava 24.57 22.58 7.73 15.73
@@
Is model size directly linked to performance? Fig. 7 illustrates the performance of a series of models with varying parameter scales. While larger models generally tend to achieve better performance, the overall performance is more influenced by the model’s architecture rather than its parameter size. For instance, the LLaVA-NeXT model with 34B parameters performs worse than the InternVL2 model with 26B parameters. Furthermore, when the parameter size of LLaVA-NeXT increases from 8B to 13B, its accuracy even drops by 0.84%.
Does GPT-4o exceed human performance? As shown in Table 2, GPT-4o’s overall accuracy surpasses human performance by 9.99%. We attribute this to the fact that our human participants are primarily students, who may lack the specialized knowledge included in GPT-4o’s extensive knowledge base and GPT-4o is the most SOTA model, which demonstrates a substantial lead over all other evaluated models. However, despite outperforming humans in most tasks, GPT-4o falls far behind in tasks such as BD and CO, which exposes its limitations in spatial perception and counting ability in agriculture. More results can be found in Appendix C.5.